A 2.5 g sample of KClO3was heated. The liberated oxygen gas was collected over water at 24 degree Celsius and 755 millimeter Hg. What was the mass and volume of the oxygen gas collected in the set up?



 2KClO 3 → 2KCl   +   3 O2 


1
Expert's answer
2021-02-01T03:43:03-0500

2KClO3 → 2KCl + 3 O2

So,

2 moles of KClO3 produce 3 moles O2

Molar mass of KClO3 = 122.55 g/mol

Molar mass of O3 = 32 g/mol


So,

(122.55×2)g KClO3 produce (32×3)g O2

2.5 g KClO3 produce = (96×2.5/245.1) g O2

= 0.9792 g O2

So, mass of oxygen gas produce = 0.9792 g


0.9792 g O2 = (0.9297/32) mol O2

So, number of moles, n = (0.9297/32) mol

Temperature, T = (24+273)K

= 297 K

Pressure,P = 755 mmHg

= (755/760) atm

R= 0.082 L.atm/(mol.K)

We know,

PV = nRT

Or, V = nRT/P

= [(0.9297/32)×0.082×297]÷(755/760)

= 0.712 L


So, mass of oxygen gas produce = 0.9792 g

volume of the produced oxygen gas = 0.712 L